# Patrick Kahardipraja1, Thomas Wiegand1,2,3, Wojciech Samek1,2,3, Sebastian Lapuschkin1

### v1

@@ Appendix A Extended related work
An output-centric approach was introduced by Gur-Arieh et al. (2025) in an attempt to address another key challenge-the feature descriptions generated via the data samples that activate the feature the most, often fail to reflect its influence on the model’s output. The study demonstrates that a combined approach, integrating both activation-based and output-based data, results in more accurate feature descriptions and improves performance in causality evaluations.
@@
Previous research consistently shows that open-source models are effective for generating explanations, with advanced models producing better descriptions. For instance, Bills et al. (2023) report that GPT-4 achieves the highest scores, whereas Claude 3.5 Sonnet performs best for Paulo et al. (2024). The number of data samples used to generate feature description also varies across studies. Bills et al. (2023) use the top five most activating samples, while Choi et al. (2024) select 10-20 of the most activating samples to generate multiple descriptions for evaluation. In contrast, Paulo et al. (2024) use 40 samples and suggests that randomly sampling from a broader set of activating leads to descriptions that cover a more diverse set of activating examples, whereas using only top activating examples often yields more concise descriptions which fails to capture the entire description.
The datasets used in these studies also differs. Paulo et al. (2024) utilize the RedPajama 10M Computer (2023) dataset, Choi et al. (2024) use the full LMSYSChat1M Zheng et al. (2023) and 10B token subset of FineWeb Penedo et al. (2024), and Bills et al. (2023) — on WebTextRadford et al. (2019) and the data used to train GPT-2 Radford et al. (2019). Additionally, different delimiter conventions are observed: Choi et al. (2024) use delimiters, Paulo et al. (2024) use << >>, and Bills et al. (2023) use numerical markers.
@@ Appendix B Data Preprocessing
For our work, we used an uncopyrighted version of the Pile dataset, with all copyrighted content removed, available on Hugging Face Gao et al. (2020) (https://huggingface.co/datasets/monology/pile-uncopyrighted). This version contains over 345.7 GB of training data from various sources. From this dataset, we extracted approximately 6 GB while preserving the relative proportions of the original data sources. The extracted portion from the training partition was used to collect the most activated samples. For evaluations, we utilized the test partition from the same dataset, applying identical preprocessing steps as those used for the training data.
@@
Our preprocessing involved several steps to ensure a balanced and informative dataset. First, we used the NLTK Bird et al. (2009) sentence tokenizer to break large text chunks into individual sentences. We then filtered out sentences in the bottom and top fifth percentiles based on length, as these were typically out-of-distribution cases consisting of single words or characters or a few outliers. This step helped achieve a more balanced distribution. Additionally, we removed sentences containing only numbers or special characters with no meaningful content. Finally, duplicate sentences were deleted.
Training Test
@@ C.1 Implementation of Automated Interpretability
Model Layers HuffingFace Descriptions
@@
Our experiments are based on models and feature descriptions presented in Table 11. We generate our feature descriptions as follows. In the implementation of the autointerpretability pipeline, we’re closely following Bills et al. (2023); Paulo et al. (2024); Rajamanoharan et al. (2024) and others: we pass the natural dataset, used for generating descriptions (see Appendix B) through the model, and via forward hooks we access the activations of each feature and each layer. This part can also be easily parallelized. With Gemma Scope SAEs, a wrapper class is implemented, that is built into the model as another module. It can be extended easily to other SAEs.
After passing the whole dataset through the model, for each feature we take top $10^{3}$ data samples based on the maximum activation of tokens. We only consider the maximum activating token of a data sample when we do the sorting. Later, we uniformly subsample the necessary number of data samples, i.e. 5, 15, …, 50, that are further passed to an explainer LLM.
The reason why we subsample from top $10^{3}$ is to avoid outliers in terms of activation. Top $10^{3}$ still represents top 0.001% of the dataset. More complex sampling strategies can bring better performance, as described in Appendix A, but their implementation and evaluation is left out for the future work.
Experiments are performed on the following models: Gemma-2-2b layer 20 Riviere et al. (2024), Llama-3.1-8B-Instruct layer 19 Dubey et al. (2024), and SAEs Gemma Scope 16K and 65K layer 20 Lieberum et al. (2024). We also generate descriptions using baseline methods, namely TF-IDF and unembedding matrix projection. Term Frequency-Inverse Document Frequency (TF-IDF) is a widely used technique in NLP for measuring the importance of a word in a document relative to a corpus. It balances word frequency with how uniquely the word appears across documents, assigning higher scores to informative words while down-weighting common ones. We generate these values using 15 maximally activating samples. On the other hand, the unembedding matrix ($W_{U}$) Joseph Bloom (2024) in transformer models maps the residual stream activations to vocabulary logits, determining word probabilities in the output. By analyzing projections onto this unembedding matrix, we gain insight into how learned features influence token predictions. To generate SAE unembedding descriptions, we generate logit weight distribution across the vocabulary, and then use the top ten words with the highest probabilities as feature description.
@@
here, $W_{U}$ is the unembedding matrix of a transformer model and $W_{dec}$ are the decoder weights of sparse auto encoders.
This reveals which words are most associated with a given feature, enabling interpretability of sparse autoencoder (SAE) features, as well as MLP neurons. Both these methods are cheap baselines to compare with different auto-interpretability generated descriptions.
C.2 Prompts Engineering
Our feature description pipeline consists of several key components. The Subject Model for which the descriptions are getting generated, while the Explainer Model is a larger model used to generate descriptions. The System Prompt provides task-specific instructions to the LLM, detailing what to focus on in the provided samples and how to format the output. We append 5, 10, or 50 sentences along with a user_message_ending at the end, which helps reinforce the expected output structure. Before normalizing activations, we first average activations for tokens belonging to the same word. These values are then normalized between 0 and 10. For delimiters, we use single curly brackets if the activation intensity is below 4 and double curly brackets otherwise. For numerical input, we provide the LLM with dictionary of most activated token at the end of each sample.
@@ C.2 Prompts Engineering
Table 13: Average number of tokens as the number of sentences increases. These values are based on tokenizer used for Gemma-Scope-16k and are not the number of tokens per requests generated by openai.
Main prompt: The main prompt for our we use to generate descriptions is given below:
@@
Prompt variation 1: using numbers, instead of delimiters for further experiments as mentioned in , we include small variations in the main prompt
@@
Prompt variation 2: main prompt + no shotsc
Prompt variation 3: main prompt + one shot example
Prompt variation 4: main prompt + five shots
@@ C.3 Computational Costs
Due to the differences in the implementation, calculating maximum activating samples is significantly cheaper for the complete models, since we are doing it at the same moment for the whole model, and the price is divided on the total number of features that we consider. For SAEs, in the current implementation we only do it per one layer, which is significantly reducing the number of considered features. In the future we are planning on optimizing this process in a way, that it is possible to calculate it at the same time for multiple chosen SAEs. The size of SAEs is also playing an important role: in terms of total costs it is more expensive, but if we consider the cost per $10^{3}$ features, the larger the SAE - the cheaper it is.
@@
Our framework is designed to work with a wide variety of subject models, including most HuggingFace Transformers (Wolf et al., 2020) as well as any feature implemented as a named module in PyTorch. Moreover, our framework is extensible: interpretability tools such as SAEs or various supervised interpretability techniques can be integrated with minimal effort, provided they implement some basic steering functions. In addition, we offer an interface for a diverse set of evaluating LLMs, whether open-weight or proprietary, with support for platforms such as vllm, Ollama, OpenAI, and Azure APIs. Our implentation can be found at https://github.com/brunibrun/FADE.
@@ D.1 Implementation and Computational Efficiency
[…]
@@
Our method allows users to control the cost by setting the number of synthetic samples (denoted $n$) relative to the full size of the natural dataset ($N$). By pre-computing activations from the natural dataset in parallel, we effectively reduce the per-run complexity from $\mathcal{O}(N*M)$ for $M$ neurons to $\mathcal{O}(n*M)$. Given that $n$ is typically in the hundreds while $N$ is in the millions, this strategy yields big efficiency gain.
Additionally, we only execute the computationally intensive faithfulness evaluation when both clarity and responsiveness exceed a user-configurable threshold. This conditional execution ensures that unnecessary computations are avoided for features that do not meet our interpretability criteria.
@@ D.2 Details on the Experiment Setup
Licenses Gemma-2-2b is released under a custom Gemma Terms of Use. Gemma Scope SAEs are released under Creative Commons Attribution 4.0 International. Llama3.1-8B-Instruct is released under a custom Llama 3.1 Community License. Transluce feature descriptions, Pile Uncopyrighted dataset and LangChain are released under MIT License. vLLM is released under Apache 2.0 License.
@@ D.2.2 Associated Cost
The activation generation for the subject models for the results in section 4.1 was run locally on a cluster with NVIDIA A100 GPUs with 40GB of VRAM. Similar to section C.3 we assume a price of \$1/hr. Since activations can be cached in parallel for the whole model, only a single pass over the evaluation dataset was needed per model. We estimate a needed time of 24 hours on one GPU for the activation generation, resulting in a cost of 24\$ per model. During the evaluations, only the activations of synthetic samples need to be computed. Due to a suboptimal configuration of the evaluation and subject LLM in our experiments we assume an average time of one minute per neuron evaluation, leading to an average cost of 0.024\$ per evaluated feature. The estimated cost for the evaluating LLM consists of the cost for the generation of synthetic samples as well as the cost for rating natural data. In the configuration used for the experiments, unless stated otherwise, we estimate the evaluating LLM creates about 2,000 tokens per feature evaluation, which corresponds to a cost of about 0.0012\$ per feature at an output token cost of 0.6\$ per million tokens. For the rating part the evaluating LLM recieves about 20,000 tokens as input, which corresponds to a cost of 0.003\$ per feature at an input token cost of 0.15\$ per million tokens. Corresponding results are presented in Table 15.
# of features evaluated GPU Time (h) Cost/ $10^{3}$ features(\$)
Gemma-2-2b 9,000 216 + 24 26.67
Gemma Scope 16K 25,000 600 + 24 24.96
Gemma Scope 65K 2,000 48 + 24 36
Llama-3.1-8B 2,000 48 + 24 36
Table 15: Computational cost comparison for running the evaluation experiments on GPU.
@@ E.1 Quantitative Analysis
SAEs concept narrowness results in lower interpretability. This might seem counterintuitive at first, but again it is important to stress that we are not evaluating the features by themselves, but instead the adequacy of the proposed feature description to these features. One potential explanation is that larger SAEs distribute concepts more finely across features. As a result, a slightly inaccurate description that might have still activated a feature in a smaller SAE may fail to activate the corresponding feature in a larger SAE, where concepts are encoded even more sparsely. The consistency of this result is demonstrated by using different feature descriptions — MaxAct*, produced in this work, and the ones available on Neuronpedia.
@@
Some Gemma-2 layers look almost not interpretable. The feature distribution in layer 12 differs significantly from other layers. Unlike other layers, layer 12 lacks the characteristic rightward elevation in clarity, responsiveness, and purity scores, as presented on Figure 6. Manual analysis of the heatmaps supported the theory, that layer 12 demonstrates a high level of polysemanticity. Interestingly, similar result was demonstrated on Llama-3.1-8B-Instruct by Choi et al. (2024). Evaluating several complete models with [Uncaptioned image]FADE would provide more insights into interpretability of different models components.
Numeric input shows marginally better performance. As illustrated in Figure 7 (a), the performance difference between numeric input-based prompts and delimiter-based highlighting is not statistically significant for both MLP neurons and SAEs, though the mean score is slightly higher for numeric input. A more comprehensive evaluation across multiple layers and a larger feature set is needed to determine the optimal approach. Nonetheless, these findings challenge prior assertions that highlighting the most activating tokens is superior due to LLMs’ assumed difficulty in processing numerical inputs Choi et al. (2024).
Increasing sample count improves performance. A clear trend emerges: providing more samples to the explainer LLM enhances its performance. However, this also increases the computational cost of feature generation due to the higher token count. While 15 samples yield strong results, 50 samples perform even better, as shown in Figure 7 (b).
@@
Increasing examples improves performance but raises costs. As shown in Figure 7 (c), increasing the number of shots consistently enhances performance on activation-based metrics without significantly affecting faithfulness, contradicting findings in Choi et al. (2024). This discrepancy may stem from differences in provided examples or feature types, as our analysis focuses on Gemma Scope SAEs, whereas Choi et al. (2024) examined Llama-3.1-8B-Instruct neurons. Additionally, a higher shot count increases computational costs due to the larger token input (see Table 12). In this study, we primarily use 2-shot prompts, balancing performance and cost efficiency.
Stronger explainer LLMs yield better feature descriptions. More capable models consistently achieve higher performance across nearly all metrics, as presented in Figure 8. Llama-3.1-70B-Instruct (quantized) performs comparably to GPT-4o mini, aligning with findings from the evaluation of LLMs (see Table 1). Performance generally declines with model size, except for Llama-3.2-1B-Instruct, which frequently fails to adhere to the required output format, leading to significantly poorer results across all metrics.
MaxAct* demonstrates superior performance. Our findings highlight the importance of considering all four [Uncaptioned image]FADE metrics when optimizing automated interpretability approaches. On Gemma Scope 16K, MaxAct* outperforms all baselines across all metrics except faithfulness (see Figure 9). We show that our generated descriptions significantly surpass those currently available on Neuronpedia. Additionally, we compare MaxAct* to TF-IDF and unembedding-based baselines (see Appendix C.1). While the unembedding method underperforms in activation-based metrics such as clarity, responsiveness, and purity, it achieves notably higher faithfulness by explicitly considering the feature’s output behavior. This underscores that faithfulness depends on both the feature type (e.g., SAEs vs. MLP neurons) and the generated description.
Figure 10: Examples of feature descriptions, obtained via MaxAct*, demonstrating low clarity, but medium or high responsiveness and purity. Feature 475: “Frequent presence of token ’self’ indicating object oriented programming concept”; feature 653: “Word ’make’ in different forms, expressions and concepts”; feature 821: “Expressions indicating lists or explanations”.

### v2

@@ Appendix A Extended related work
An output-centric approach was introduced by Gur-Arieh et al. (2025) in an attempt to address another key challenge – the feature descriptions generated via the data samples that activate the feature the most, often fail to reflect its influence on the model’s output. The study demonstrates that a combined approach, integrating both activation-based and output-based data, results in more accurate feature descriptions and improves performance in causality evaluations.
@@
Previous research consistently shows that open-source models are effective for generating explanations, with advanced models producing better descriptions. For instance, Bills et al. (2023) report that GPT-4 achieves the highest scores, whereas Claude 3.5 Sonnet performs best for Paulo et al. (2024). The number of data samples used to generate feature descriptions also vary across studies. Bills et al. (2023) use the top five most activating samples, while Choi et al. (2024) select 10-20 of the most activating samples to generate multiple descriptions for evaluation. In contrast, Paulo et al. (2024) use 40 samples and suggest that randomly sampling from a broader set of activating leads to descriptions that cover a more diverse set of activating examples, whereas using only top activating examples often yields more concise descriptions that fail to capture the entire description.
The datasets used in these studies also differ. Paulo et al. (2024) utilize the RedPajama 10M Computer (2023) dataset, Choi et al. (2024) use the full LMSYSChat1M Zheng et al. (2023) and 10B token subset of FineWeb Penedo et al. (2024), Bills et al. (2023) – WebTextRadford et al. (2019) and the data used to train GPT-2 Radford et al. (2019). Additionally, different delimiter conventions are observed: Choi et al. (2024) use delimiters, Paulo et al. (2024) use << >>, and Bills et al. (2023) use numerical markers.
@@ Appendix B Data Preprocessing
For our work, we used an uncopyrighted version of the Pile dataset, with all copyrighted content removed, available on Hugging Face Gao et al. (2020) (https://huggingface.co/datasets/monology/pile-uncopyrighted). This version contains over 345.7 GB of training data from various sources. From this dataset, we extracted approximately 6 GB for labelling while preserving the relative proportions of the original data sources. The extracted portion from the training partition was used to collect the most activated samples. For evaluations, we utilized the test partition from the same dataset, applying identical preprocessing steps as those used for the training data.
@@
Our preprocessing involved several steps to ensure a balanced and informative dataset. First, we used the NLTK Bird et al. (2009) sentence tokenizer to split large text chunks into individual sentences. We then filtered out sentences in the bottom and top fifth percentiles based on length, as these were typically out-of-distribution cases consisting of single words, characters, or a few outliers. This step helped achieve a more balanced distribution. Additionally, we removed sentences containing only numbers or special characters with no meaningful content. Finally, duplicate sentences were deleted.
Labeling Dataset Evaluation Dataset
@@ C.1 Implementation of Automated Interpretability
Model Layers HuggingFace Descriptions
@@
Our experiments are based on models and feature descriptions presented in Table 11. We generate our feature descriptions as follows. In the implementation of the autointerpretability pipeline, we are closely following Bills et al. (2023); Paulo et al. (2024); Rajamanoharan et al. (2024) and others: we pass the natural dataset, used for generating descriptions (see Appendix B) through the model, and via forward hooks we access the activations of each feature and each layer, which can be done in parallel. For Gemma Scope SAEs we implement a wrapper class that is built into the model as named module, and can be easily extended to other SAEs.
After passing the whole dataset through the model, we take top $10^{3}$ data samples for each feature, based on the maximum activation of tokens. We only consider a single maximum activating token of a data sample. Later, we uniformly subsample the necessary number of data samples, i.e. 5, 15, …, 50, that are further passed to an explainer LLM. We do that to avoid outliers – a small part of the data which activates the feature the most, but does not represent its overall behavior well. The considered top $10^{3}$ most activating samples represent top 0.001% of the dataset. More complex sampling strategies can yield better performance, as described in Appendix A, but their implementation and evaluation is left out for future work.
Experiments are performed on the following models: Gemma-2-2b layer 20 Riviere et al. (2024), Llama-3.1-8B-Instruct layer 19 Dubey et al. (2024), and SAEs Gemma Scope 16K and 65K layer 20 Lieberum et al. (2024). In addition, we generate descriptions using baseline methods, namely TF-IDF and unembedding matrix projection. Term Frequency-Inverse Document Frequency (TF-IDF) is a widely used technique in NLP for measuring the importance of a word in a document relative to a corpus. It balances word frequency with how uniquely the word appears across documents, assigning higher scores to informative words while down-weighting common ones. We generate these values using 15 maximally activating samples. On the other hand, the unembedding matrix ($W_{U}$) Joseph Bloom (2024) in transformer models maps the residual stream activations to vocabulary logits, determining word probabilities in the output. By analyzing projections onto this unembedding matrix, we gain insight into how learned features influence token predictions. To generate SAE unembedding descriptions, we generate a logit weight distribution across the vocabulary, and then use the top ten words with the highest probabilities as feature descriptions.
@@
where $W_{U}$ is the unembedding matrix of a transformer model and $W_{dec}$ are the decoder weights of sparse auto encoders.
This reveals which words are most associated with a given feature, enabling interpretability of sparse autoencoder (SAE) features, as well as MLP neurons. Both these methods are cheap baselines to compare with descriptions generated via different autointerpretability approaches.
C.2 Prompt Engineering
Our feature description pipeline consists of several key components, such as the Subject Model, for which the descriptions are generated, the Explainer Model, which is a larger language model, used to generate descriptions, and the System Prompt, which provides task-specific instructions to the Explainer Model, detailing what to focus on in the provided samples and how to format the output. We append 5, 10, or 50 sentences along with a user_message_ending at the end, which helps reinforce the expected output structure. Before normalizing activations, we first average activations for tokens belonging to the same word. These values are then normalized between 0 and 10. For delimiters, we use single curly brackets if the activation intensity is below 4, and double curly brackets otherwise. For numerical input we provide the Explainer Model with a dictionary of the most activated tokens after each sample.
@@ C.2 Prompt Engineering
Table 13: Average number of tokens as the number of sentences increases. These values are based on tokenizer used for Gemma-Scope-16k and are not the number of tokens per request generated by OpenAI.
Main prompt:
The main prompt we use to generate descriptions is given below.
@@
Prompt variation 1:
Main prompt + dictionary of most activated tokens with their normalized activation values.
@@
Prompt variation 2: Main prompt + zero shots.
Prompt variation 3: Main prompt + one shot.
Prompt variation 4: Main prompt + five shots.
@@ C.3 Computational Costs
Due to the differences in the implementation, calculating maximum activating samples is significantly cheaper for the complete models, since we do it during the same forward pass for the whole model, and the price is divided by the total number of features we consider. For SAEs, in the current implementation we only do it per one layer, which significantly reduces the number of considered features. In the future we are planning on optimizing this process in a way, that it is possible to calculate it simultaneously for multiple chosen SAEs. The size of SAEs also plays an important role: in terms of total costs it is more expensive, but if we consider the cost per $10^{3}$ features, the larger the SAE – the cheaper it is.
@@ D.1 Implementation and Computational Efficiency
To compute the Purity and Responsiveness, we sample evaluation sequences from the natural dataset based on the activation distribution of the considered feature. For each sequence, we calculate the maximum absolute activation value across all tokens. Using these values, we sample sequences by selecting a user-configurable percentage of those with the strongest activations and for the remainder, drawing an equal number of samples from each of the following percentile ranges: $[0\%,50\%)$, $[50\%,75\%)$, $[75\%,95\%)$, and $[95\%,100\%]$.
@@
Our method allows users to control the cost by setting the number of synthetic samples (denoted $n$) relative to the full size of the natural dataset ($N$). By pre-computing activations from the natural dataset in parallel, we effectively reduce the per-run complexity from $\mathcal{O}(N*M)$ for $M$ neurons to $\mathcal{O}(n*M)$. Given that $n$ is typically in the hundreds while $N$ is in the millions, this strategy yields significant efficiency gains.
Additionally, we only execute the computationally costly Faithfulness evaluation when both Clarity and Responsiveness exceed a user-configurable threshold. This conditional execution ensures that unnecessary computations are avoided for features that do not meet our interpretability criteria.
@@ D.2 Details on the Experiment Setup
Licenses
Gemma-2-2b is released under a custom Gemma Terms of Use. Gemma Scope SAEs are released under Creative Commons Attribution 4.0 International. Llama3.1-8B-Instruct is released under a custom Llama 3.1 Community License. Transluce feature descriptions, Pile Uncopyrighted dataset and LangChain are released under MIT License. vLLM is released under Apache 2.0 License.
@@ D.2.2 Associated Cost
The activation generation for the Subject Models for the results in section 4.1 was run locally on a cluster with NVIDIA A100 GPUs with 40GB of VRAM. Similar to section C.3 we assume a price of \$ 1/hr. Since activations can be cached in parallel for the whole model, only a single pass over the evaluation dataset was needed per model. We estimate a needed time of 24 hours on one GPU for the activation generation, resulting in a cost of \$24 per model. During the evaluations, only the activations of synthetic samples and Faithfulness text generations need to be computed. Given our hyper-parameter setting described in D.2, we measure an average of 31 seconds GPU time per neuron. Of that, roughly 2 seconds are spent on the Clarity, Responsiveness and Purity metrics, while the much more expensive Faithfulness computation contributes the remaining 29 seconds. These 29 seconds arise because, with our present threshold of 0.5, the Faithfulness experiment is executed on 53% of candidate labels; if it were executed every time it would add roughly 55 seconds. This leads us to an average cost of \$0.0086 per neuron, excluding the fixed activation caching cost.
@@
The estimated cost for the evaluating LLM consists of the cost for the generation of synthetic samples as well as the cost for rating natural data. In the configuration used for the experiments, unless stated otherwise, we estimate the evaluating LLM creates on average about 2800 tokens per feature evaluation, which corresponds to a cost of about \$0.00168 per feature at an output token cost of \$0.6 per million tokens. For the rating part, the evaluating LLM receives about 23000 tokens as input, of which on average 17000 are from the Responsiveness and Purity experiments and 6000 are from the Faithfulness experiments, which corresponds to a total cost of \$0.00345 per feature at an input token cost of \$0.15 per million tokens. Corresponding results are presented in Table 15.
Features evaluated GPU Time (h) Inference Cost (\$) Cost per feature (\$)
Gemma-2-2b 9.000 24+77.5 46.17 0.0164
Gemma Scope 16K 25.000 24+215.3 128.25 0.0147
Gemma Scope 65K 2.000 24+17.2 10.26 0.0257
Llama-3.1-8B 2.000 24+17.2 10.26 0.0257
Table 15: Computational cost estimates of the evaluation experiments.
The total cost in our experiments is primarily driven by the use of high values of hyperparameters intended to ensure high evaluation accuracy as outlined in section D.2. Here especially the low threshold of executing the Faithfulness experiments is a big contributor to overall cost. Increasing this threshold will reduce the computational cost substantially. Additionally, the high number of samples that is being rated for the activation-based metrics can be reduced in order to significantly lower the amount of needed inference tokens for the evaluating LLM. Furthermore, as seen in Table 1, often the size of the evaluating LLM itself can be reduced with minimal accuracy losses, resulting in cheaper inference costs via e.g. self-hosted models or simply cheaper inference. For example if we had run the Faithfulness experiments only on the most promising 25% percent of labels, corresponding to a Faithfulness threshold of approximately 0.9 given our evaluated labels, the average GPU compute cost per neuron, excluding the static computation of cached activations, would decrease from \$0.0086 to \$0.0044.
In our experiments we chose these high hyperparameter settings, as we wanted to prioritize robustness and replicability of our results. We believe however, that even with significantly lower configurations, especially regarding the Faithfulness-threshold and number of modification factors, valuable insights can be found at a considerably lower cost.
D.3 Combining [Uncaptioned image]FADE metrics into a single score
The principal value of [Uncaptioned image]FADE lies in its multidimensional analysis of feature-to-description alignment. Reducing these dimensions to a single score may obscure important nuances that are crucial for developing and refining interpretability techniques. Separate metrics allow researchers to pinpoint specific weaknesses in their descriptions and address them directly. However, for applications that require simplified comparison or ranking of descriptions, appropriate combination methods can be selected based on the specific requirements of the use case. We propose that various means (weighted, geometric, or harmonic) could serve the purpose of creating one holistic evaluation metric. A weighted combination of the four metrics would allow one to adjust importance based on specific use cases and applications. The geometric mean would be particularly suitable, as it accounts for all metrics while being naturally sensitive to underperforming dimensions. Alternatively, a harmonic mean could be employed for cases where balanced performance is even more critical, as it penalizes low scores more severely than the geometric mean.
@@ E.1 Quantitative Analysis
SAEs concept narrowness results in lower interpretability
This might seem counterintuitive at first, but it is important to stress again that we are not evaluating the features by themselves, but instead the adequacy of the proposed feature description to these features. One potential explanation is that larger SAEs distribute concepts more finely across features. As a result, a slightly inaccurate description that might have still activated a feature in a smaller SAE may fail to activate the corresponding feature in a larger SAE, where concepts are encoded even more sparsely. The consistency of this result is demonstrated by using different feature descriptions – MaxAct*, produced in this work, and the ones available on Neuronpedia.
@@
Some Gemma-2 layers look almost not interpretable
The feature distribution in layer 12 differs significantly from other layers, since it lacks the characteristic rightward elevation in Clarity, Responsiveness, and Purity scores, as presented on Figure 7. Manual analysis of the heatmaps support the theory, that layer 12 demonstrates a high level of polysemanticity. Interestingly, similar result were demonstrated on Llama-3.1-8B-Instruct by Choi et al. (2024). Evaluating several complete models with [Uncaptioned image]FADE would provide more insights into the interpretability of different models components.
Numeric input shows marginally better performance
As illustrated in Figure 8 (a), the performance difference between numeric input-based prompts and delimiter-based highlighting is not statistically significant for both MLP neurons and SAEs, though the mean score is slightly higher for numeric input. A more comprehensive evaluation across multiple layers and a larger feature set is needed to determine the optimal approach. Nonetheless, these findings challenge prior assertions that highlighting the most activating tokens is superior due to LLMs’ assumed difficulty in processing numerical inputs Choi et al. (2024).
Increasing sample count improves performance
A clear trend emerges: providing more samples to the explainer LLM enhances its performance. However, this also increases the computational cost of feature generation due to the higher token count. While 15 samples yield strong results, 50 samples perform even better, as shown in Figure 8 (b).
@@
Increasing examples improves performance but raises costs
As shown in Figure 8 (c), increasing the number of shots consistently enhances performance on activation-based metrics without significantly affecting Faithfulness, contradicting findings in Choi et al. (2024). This discrepancy may stem from differences in provided examples or feature types, as our analysis focuses on Gemma Scope SAEs, whereas Choi et al. (2024) examined Llama-3.1-8B-Instruct neurons. Additionally, a higher shot count increases computational costs due to the larger token input (see Table 12). In this study, we primarily use 2-shot prompts, balancing performance and cost efficiency.
Stronger explainer LLMs yield better feature descriptions
More capable models consistently achieve higher performance across nearly all metrics, as presented in Figure 9. Llama-3.1-70B-Instruct (quantized) performs comparably to GPT-4o mini, aligning with findings from the evaluation of LLMs (see Table 1). Performance generally declines with model size, except for Llama-3.2-1B-Instruct, which frequently fails to adhere to the required output format, leading to significantly poorer results across all metrics.
MaxAct* demonstrates superior performance
Our findings highlight the importance of considering all four [Uncaptioned image]FADE metrics when optimizing automated interpretability approaches. On Gemma Scope 16K, MaxAct* outperforms all baselines across all metrics except Faithfulness (see Figure 10). We show that our generated descriptions significantly surpass those currently available on Neuronpedia. Additionally, we compare MaxAct* to TF-IDF and unembedding-based baselines (see Appendix C.1). While the unembedding method underperforms in activation-based metrics such as Clarity, Responsiveness, and Purity, it achieves notably higher Faithfulness by explicitly considering the feature’s output behavior. This underscores that Faithfulness depends on both the feature type (e.g., SAEs vs. MLP neurons) and the generated description.
MaxAct* 16K MaxAct* 65K Neuronpedia 16K Neuronpedia 65K
C R P F C R P F C R P F C R P F
Clarity 1.00 0.68 0.63 0.17 1.00 0.69 0.65 0.14 1.00 0.63 0.53 0.21 1.00 0.56 0.49 0.28
Responsiveness - 1.00 0.86 0.16 - 1.00 0.92 0.10 - 1.00 0.88 0.21 - 1.00 0.87 0.14
Purity - - 1.00 0.12 - - 1.00 0.12 - - 1.00 0.24 - - 1.00 0.23
Faithfulness - - - 1.00 - - - 1.00 - - - 1.00 - - - 1.00
Table 16: Correlation of FADE metrics for different sets of feature descriptions.